# NJL-Chiral Soliton and the Nucleon Equation of State at supra-saturation density: Impact of Chiral Symmetry Restoration

### v1

@@ Abstract
It has been conjectured that, at sufficiently high baryon densities, the equation of state (EoS) of bulk nuclear matter can be identified with that of the nucleon core. In this work, we illustrate how the energy density and pressure distributions inside individual nucleons can be utilized to construct the EoS of supra-dense matter. In our framework, nucleons arise as topological solitons stabilized by vector mesons, which are dynamically generated through the path integral bosonization of an underlying Nambu–Jona-Lasinio (NJL) model. The restoration of chiral symmetry is implemented dynamically via a self-consistent, density-dependent scalar field, which modifies the (isovector) and (isoscalar) channels of the soliton. We analyze the resulting changes in soliton properties for different NJL parameter sets and demonstrate that the progressive restoration of chiral symmetry leads to a stiffening of the soliton-based EoS, making it compatible with existing neutron star EoSs. An EoS constructed from the solutions of the energy-density and pressure profiles at the center of the nucleon is also explored.
NJL Model, Nucleon Structure, Chiral Soliton, quark matter
I Introduction
@@
The phase transition of Quantum Chromodynamics (QCD) from nuclear matter into a quark-gluon plasma (QGP) is generally accompanied by chiral symmetry restoration and quark deconfinement [1]. Theoretically, due to asymptotic freedom, perturbative QCD calculations are only applicable at short distances or high energy scales [2, 3, 4, 5]. At low densities and high temperatures, results from lattice QCD simulations indicate a smooth variation in thermodynamic properties, such as pressure ($p$), energy density ($\varepsilon$), and entropy with temperature. In contrast, phenomenological approaches like the Hadron Resonance Gas (HRG) model, which treats hadrons as non-interacting point particles (ideal gas) with the Hagedorn density of states [6, 7, 8, 9], predict a singular (blow-up) behavior in thermodynamic quantities near the Hagedorn temperature (chemical freeze-out). The singularity behavior is further tamed by excluded volume corrections that consider the finite hadron size, enabling a smooth crossover to quark matter [10], which suggests that the inter-hadron interactions play a crucial role. At high temperatures and densities, however, strong nucleon-nucleon interactions invalidate the HRG’s independent-particle picture, and as quark-gluon dynamics gradually dominate, the boundary between hadronic and partonic phases blurs. Hence, a theoretical unified description is demanding to account both hadronic and partonic degrees of freedom (d.o.f) in-order to describe the thermodynamic properties of dense matter.
@@
Though lattice QCD has significantly advanced our understanding of the equation of state (EoS) or the nature of the phase transition at low density and high temperature regime [11, 12, 13, 14, 15, 16], it encounters the sign problem in the high density and low temperature regime [17]. A promising avenue to understand the matter at high densities lies in the study of astrophysical systems, in particular neutron stars (NSs). The densities realized in the inner cores of massive NSs exceed those achievable in terrestrial laboratories, making NSs unique natural laboratories for probing strongly interacting matter at extreme baryon densities (supra-dense matter). Although QCD cannot yet be solved directly at the relevant density regimes, the NS interior can be modeled through EoS based on microscopic many-body calculations or phenomenological approaches. These EoSs establish a connection between the microphysics of dense matter and global NS observables such as mass, radius, tidal deformability, and moment of inertia. Observational measurements have significantly tightened constraints on the NS EoS. In particular, precise measurements of the massive $\sim 2M_{\odot}$ NS mass  [18, 19, 20, 21] and the joint mass–radius measurements from the NICER mission [22, 23, 24, 25] have placed strong constraints on the range of viable models. Moreover, the ground-breaking detection of gravitational waves (GWs) from the binary NS merger event GW170817 [26, 27, 28, 29, 30], with its measurement of tidal deformability and NSs masses, has opened a new era of multi-messenger astronomy, providing complementary constraints on the dense matter EoS [31]. Combined inputs from electromagnetic and GW observations have been widely employed to reduce EoS uncertainties and to explore the possible existence and nature of a hadron–quark phase transition in the interior of NSs [32, 33, 34, 35, 36, 37, 38, 39]. Moreover, it has been demonstrated that simultaneously incorporating the astrophysical constraints and/or the low-density constraints from chiral effective field theory (EFT) together with high-density limits from perturbative QCD provides information on the phase transition and novel constraints on the NS EoS [40, 41, 42, 43, 44, 45], with the limitations expressed in Refs. [46, 47].
@@
What if supra-dense matter already exists in finite-size nuclei? Nucleons, for instance, support densities that far exceed nuclear saturation density. In such a case, one should find a way to extract the thermodynamic properties of matter at these extreme densities. The low-energy structure and dynamics of hadrons, including nucleons and pions, are governed by spontaneously broken chiral symmetry. Models based on chiral symmetry often view nucleons as a compact hard core that contains the three valence quarks and a surrounding quark–antiquark (pion) cloud. Furthermore, the idea of a compact core in the center of nucleons has also been supported by deep-inelastic scattering data as well as from the evaluation of nucleon size by high-energy nucleus-nucleus cross sections [48, 49, 50]. The representation of the nucleon as the compact hard core with a surrounding meson cloud also serves as the key idea for the cloudy MIT bag model  [51, 52] and the chiral soliton models  [53, 54, 55, 56, 57, 58, 59]. One central idea of this paper, following the work of Ref. [58], is to assume that at sufficiently high baryonic density, the thermodynamic pressure of bulk matter can be identified with the mechanical pressure in the hard core, in short, to identify the EoS of the nucleon core with the EoS of bulk matter.
@@
Since a rigorous derivation from QCD remains elusive, it is conjectured that baryons emerge as solitons. The Skyrme model, proposed initially by [60], describes baryons as topological solitons (Skyrmions) of the nonlinear pion field. While the model’s kinetic term alone cannot support stable solutions due to Derrick’s theorem [61], stability is achieved through the addition of the quartic term (Skyrme term). A significant refinement came with the inclusion of vector mesons, motivated by the Vector Meson Dominance (VMD) framework and the central role of the $\rho$ and $\omega$ mesons in $\pi\pi$, $\pi N$, and nucleon–nucleon interactions. In Ref. [53], it has been shown that coupling the $\omega$ meson alone can stabilize solitons without requiring the Skyrme term. At the same time, the inclusion of the $\rho$ meson has improved the description of the chiral soliton [54]. The combined inclusion (axial) vector mesons $\omega$, $a$, and $\rho$ mesons, as discussed in Refs. [55, 56, 57], provide a more comprehensive description of soliton dynamics and baryon properties. These vector meson extensions enhance the model’s predictive power for nucleon structure, including electromagnetic form factors and baryon mass spectra, aligning closely with experimental observations.
@@
Numerous efforts have been devoted to derive effective hadron (nucleon and mesons) theories from the Nambu–Jona-Lasinio (NJL) model [55, 62, 59, 63, 64, 65, 66, 67, 68]. These studies demonstrate that, at low energies, the structure of the effective meson Lagrangian is largely dominated by chiral symmetry. Bosonization of the NJL model yields an effective mesonic theory that reproduces the low-energy phenomenology of pseudoscalar (pion) and (axial-)vector mesons with reasonable accuracy. In addition, self-consistent soliton solutions of the bosonized NJL model have been extensively investigated to describe nucleons. Various approaches have been developed to describe these solitons and to connect nucleon properties with experimental data, including the emergence of chiral solitons from the bosonized NJL model [69, 70, 71, 59], quark–diquark models [72, 64, 65], quark–meson descriptions, and structure function methods [73, 74], as well as the appearance of $N$ and $N^{*}$ states in the parity-doublet model. Such soliton-based models provide a transparent mechanism by which baryons can be understood as solitons of effective meson theories rooted in an underlying quark-level description. These NJL chiral soliton solutions can be compared with the Skyrme type model by a gradient expansion in the limit of infinitely heavy (axial-) vector mesons [75]. Further extensions have also addressed the role of strangeness, providing insights into hyperons in the context of NJL-based chiral solitons [76, 71].
@@
In-medium modifications of soliton properties, particularly in connection with the (fully or partial) restoration of chiral symmetry at finite density, have been investigated in several works, see Refs. [77, 78, 79]. In-medium effects are implemented through density-dependent values of the pion decay constant as well as the in-medium pion and sigma masses in a quark-meson theory in Ref. [77]. It has ben suggested that the r.m.s charge radius of the soliton may increase by up to 17% near nuclear saturation density $\rho_{\rm sat}\sim 0.16\rm\ fm^{-3}$, and a phase transition from the chirally broken Goldstone phase to the chiral restored phase occurs near a baryon density of $\sim 3\rho_{\rm sat}$ [77].
@@
Another analysis suggests that the soliton solution vanishes near a critical quark chemical potential, indicating a phase transition where the bound soliton state dissolves to the constituents before the restored chiral symmetry [78]. Furthermore, within a soliton bag model theory in the Wigner-Seitz approximation to find the quark wavefunction and bound sigma fields, it was predicted that an equilibrium among the nuclear and quark phase could exist around $\sim 6\rho_{\rm sat}$ and a complete quark plasma phase could be expected at $\sim 12\rho_{\rm sat}$ [80]. Additionally, the soliton properties at finite temperature and density have also been examined, see Refs. [78, 81, 82, 83]. The soliton description of baryons with vector mesons has also been employed to understand the properties of nuclear matter, such as incompressibility and binding energy, both in the Wigner-Seitz cell and the multi-skrymion approach. Since solitons are characterized by their r.m.s charge radii, this allows a self-consistent formalism of incorporating finite-size effects into nuclear and NS matter, linking the evolution of baryon size directly to the underlying soliton description [84].
@@
As discussed in Ref. [58], the core and cloud picture provides a valuable framework to interpret baryonic structure at finite density. The outer, soft multi-pion cloud is directly linked to spontaneously broken chiral symmetry and the approximate Nambu–Goldstone nature of the pion. Its spatial extent is expected to grow with increasing density, consistent with the reduction of the in-medium pion decay constant. In contrast, the inner baryonic core is dominated by gluonic dynamics and is not directly tied to chiral symmetry, making it comparatively robust against moderate changes in density until the compact cores approach each other and begin to overlap. Well before this overlap sets in, the mesonic tails generate meson-exchange forces near saturation density ($\rho_{\rm sat}$). By further increasing the density, quark–antiquark pairs gradually delocalize, enhancing their mobility across multiple nucleons in a percolation-like process. This regime has been referred to as “soft” deconfinement. At sufficiently high densities, however, when the baryonic hard cores themselves overlap, quark matter is realized unambiguously. This transition is termed “hard deconfinement” and may be identified with the deconfinement of valence quarks, at which point the properties of dense quark matter are governed primarily by quark degrees of freedom. It can be quantitatively assessed from the internal nucleon hard core substructure  [^1] In Ref. [58], the discussed nucleon EoS corresponds to the de-confined quark matter (valid near the hard core overlap $\sim 8\rho_{\rm sat}$). However the EoS is represented as “nucleon EoS” as it is derived from the core of the nucleon..
@@
In the work presented in Ref. [58], the EoS for dense quark matter in the hard-deconfined phase is derived from the local nucleon EoS associated with the soliton’s hard core, constructed as a skyrmion of the pion field, evaluated under vacuum conditions. Although the restoration of chiral symmetry at high densities represents a natural physical expectation, its effects were not explicitly incorporated in the model. Instead, a qualitative argument was proposed to rescale the energy density and pressure using the in-medium pion decay constant as an approximation to account for chiral symmetry restoration. In contrast, the present study aims for a self-consistent approach to quantify the influence of chiral symmetry restoration on the density-dependent evolution of nucleon properties and the EoS. A central motivation of our framework is the description of the nucleon as a topological soliton emerging from the chiral dynamics of the NJL model, embedded within an in-medium effective background density-dependent scalar field that encapsulates the progressive restoration of chiral symmetry. As described below, the parameters of this NJL model will be adjusted to reproduce the skyrmion solution under the vacuum condition used in Ref. [58].
@@
To achieve the aim of this work, we leverage the path-integral bosonization approach of the underlying NJL model, as developed in Refs. [85, 68]. In summary, the NJL model is bosonized to express quark-level interactions in terms of mesonic degrees of freedom, facilitating the description of hadronic properties. The effective Lagrangian of the NJL model, after bosonization, includes scalar, vector, and pseudoscalar meson fields. The formalism establishes direct connections between the NJL parameters and the vacuum and in-medium modifications to key meson properties, including the pion mass, pion decay constant, vector coupling constant, and vector meson masses, all of which are evaluated in terms of the density-dependent scalar field. This formulation enables us to have a systematic accounting of the connection among the NJL model parameters and to both vacuum and in-medium nucleon properties, such as the soliton’s mass, radii, and EoS, while facilitating the exploration of scalar-field-induced or the restored chiral symmetry effects on mesons and solitons in dense nuclear matter.
@@
This paper is organized as follows. In section II, the theoretical framework is provided, including the bosonized NJL model and the construction of the chiral soliton ansatz with vector mesons. In this section, details about the numerical implementation and its reliability by benchmarking it against existing results are presented as well. We discuss the sensitivity of the vacuum solution properties to the NJL model parameters and the effect of the chiral symmetry restoration on the nucleon structure and the resulting EoS in section III. Finally, we summarise our findings and discuss the broader implications in section IV.
II Theoretical Framework

### v2

@@ Abstract
It has been conjectured that, at sufficiently high baryon densities, the equation of state (EoS) of bulk nuclear matter can be identified with that of the nucleon core. In this work, we illustrate how the energy density and pressure distributions inside individual nucleons can be utilized to construct the EoS of supra-dense matter. In our framework, nucleons arise as topological solitons stabilized by vector mesons, which are dynamically generated through the path integral bosonization of an underlying Nambu–Jona-Lasinio (NJL) model. The restoration of chiral symmetry is implemented dynamically via a self-consistent, density dependent scalar field, which modifies the (isovector) and (isoscalar) channels of the soliton. We analyze the resulting changes in soliton properties for different NJL parameter sets and demonstrate that the progressive restoration of chiral symmetry leads to a stiffening of the soliton-based EoS, making it compatible with existing neutron star EoSs. An EoS constructed from the solutions of the energy-density and pressure profiles at the center of the nucleon is also explored.
journal: Eur. Phys. J. A
1 Introduction
@@
The phase transition of Quantum Chromodynamics (QCD) from nuclear matter into a quark-gluon plasma (QGP) is generally accompanied by chiral symmetry restoration and quark deconfinement RISCHKE2004 . Theoretically, due to asymptotic freedom, perturbative QCD calculations are only applicable at short distances or high energy scales Blaizot2003 ; Kajantie2003 ; Kurkela2010 ; Collins2011 . At low densities and high temperatures, results from lattice QCD simulations indicate a smooth variation in thermodynamic properties, such as pressure ($p$), energy density ($\varepsilon$), and entropy with temperature. In contrast, phenomenological approaches like the Hadron Resonance Gas (HRG) model, which treats hadrons as non-interacting point particles (ideal gas) with the Hagedorn density of states CABIBBO197567 ; Hagedorn:1967tlw ; Ericson:2016sai ; Broniowski2000 , predict a singular (blow-up) behavior in thermodynamic quantities near the Hagedorn temperature (chemical freeze-out). The singularity behavior is further tamed by excluded volume corrections that consider the finite hadron size, enabling a smooth crossover to quark matter Andronic , which suggests that the inter-hadron interactions play a crucial role. At high temperatures and densities, however, strong nucleon-nucleon interactions invalidate the HRG’s independent-particle picture, and as quark-gluon dynamics gradually dominate, the boundary between hadronic and partonic phases blurs. Hence, a theoretical unified description is demanding to account both hadronic and partonic degrees of freedom (d.o.f) in-order to describe the thermodynamic properties of dense matter.
@@
Though lattice QCD has significantly advanced our understanding of the equation of state (EoS) or the nature of the phase transition at low density and high temperature regime Aoki2006 ; Cheng2008 ; Bernard2004 ; Bazavov2009 ; Ueda2014 ; Guenther2020 , it encounters the sign problem in the high density and low temperature regime NAGATA2022 . A promising avenue to understand the matter at high densities lies in the study of astrophysical systems, in particular neutron stars (NSs). The densities realized in the inner cores of massive NSs exceed those achievable in terrestrial laboratories, making NSs unique natural laboratories for probing strongly interacting matter at extreme baryon densities (supra-dense matter). Although QCD cannot yet be solved directly at the relevant density regimes, the NS interior can be modeled through EoS based on microscopic many-body calculations or phenomenological approaches. These EoSs establish a connection between the microphysics of dense matter and global NS observables such as mass, radius, tidal deformability, and moment of inertia. Observational measurements have significantly tightened constraints on the NS EoS. In particular, precise measurements of the massive $\sim 2M_{\odot}$ NS mass  Demorest2010 ; Antoniadis:2013pzd ; Cromartie:2019kug ; Fonseca_2021 and the joint mass–radius measurements from the NICER mission miller_m_c_2019_3473466 ; Riley_2019 ; Miller_2021 ; Riley_2021 have placed strong constraints on the range of viable models. Moreover, the ground-breaking detection of gravitational waves (GWs) from the binary NS merger event GW170817 AbbottPRX ; Abbott_GWTC1 ; AbbottAJL848 ; AbbottPRL119 ; RICHABBOTT2021 , with its measurement of tidal deformability and NSs masses, has opened a new era of multi-messenger astronomy, providing complementary constraints on the dense matter EoS AbbottPRL121 . Combined inputs from electromagnetic and GW observations have been widely employed to reduce EoS uncertainties and to explore the possible existence and nature of a hadron–quark phase transition in the interior of NSs Most2018 ; Gloria2019 ; Somasundaram:2022 ; Roy2024 ; Christian2023a ; Christian2023b ; Pfaff2021 ; Somasundaram:2023a ; HAN2023913 . Moreover, it has been demonstrated that simultaneously incorporating the astrophysical constraints and/or the low-density constraints from chiral effective field theory (EFT) together with high-density limits from perturbative QCD provides information on the phase transition and novel constraints on the NS EoS Shirke2022 ; Annala2019 ; Annala2022 ; Annala2023 ; Komoltsev2025 ; Gorda2023 ; Fan2024 , with the limitations expressed in Refs. Somasundaram:2023b ; Komoltsev2023 .
@@
What if supra-dense matter already exists in finite-size nuclei? Nucleons, for instance, support densities that far exceed nuclear saturation density. In such a case, one should find a way to extract the thermodynamic properties of matter at these extreme densities. For instance several recent works have related the high-density environment of compact star cores to the mechanical properties of the nucleon, highlighting possible bridges between hadronic structure and NS matter Lorce2018 ; Fukushima2020 . The low-energy structure and dynamics of hadrons, including nucleons and pions, are governed by spontaneously broken chiral symmetry. Models based on chiral symmetry often view nucleons as a compact hard core that contains the three valence quarks and a surrounding quark–antiquark (pion) cloud. Furthermore, the idea of a compact core in the center of nucleons has also been supported by deep-inelastic scattering data as well as from the evaluation of nucleon size by high-energy nucleus-nucleus cross sections HERA_2006 ; Xiong2019 ; Nijs2022 . The representation of the nucleon as the compact hard core with a surrounding meson cloud also serves as the key idea for the cloudy MIT bag model  Thomas1981 ; Miller2002 and the chiral soliton models  ADKINS1984251 ; IGARASHI1985 ; Alkofer1994 ; Meissner1986 ; Meissner1987 ; Fukushima2020 ; Zuckert1997 . One central idea of this paper, following the work of Ref. Fukushima2020 , is to assume that at sufficiently high baryonic density, the thermodynamic pressure of bulk matter can be identified with the mechanical pressure in the hard core, in short, to identify the EoS of the nucleon core with the EoS of bulk matter.
@@
Since a rigorous derivation from QCD remains elusive, it is conjectured that baryons emerge as solitons. The Skyrme model, proposed initially by Skyrme1961 , describes baryons as topological solitons (Skyrmions) of the nonlinear pion field. While the model’s kinetic term alone cannot support stable solutions due to Derrick’s theorem Derrick1964 , stability is achieved through the addition of the quartic term (Skyrme term). A significant refinement came with the inclusion of vector mesons, motivated by the Vector Meson Dominance (VMD) framework and the central role of the $\rho$ and $\omega$ mesons in $\pi\pi$, $\pi N$, and nucleon–nucleon interactions. In Ref. ADKINS1984251 , it has been shown that coupling the $\omega$ meson alone can stabilize solitons without requiring the Skyrme term. At the same time, the inclusion of the $\rho$ meson has improved the description of the chiral soliton IGARASHI1985 . The combined inclusion (axial) vector mesons $\omega$, $a$, and $\rho$ mesons, as discussed in Refs. Alkofer1994 ; Meissner1986 ; Meissner1987 , provide a more comprehensive description of soliton dynamics and baryon properties. These vector meson extensions enhance the model’s predictive power for nucleon structure, including electromagnetic form factors and baryon mass spectra, aligning closely with experimental observations.
@@
Numerous efforts have been devoted to derive effective hadron (nucleon and mesons) theories from the Nambu–Jona-Lasinio (NJL) model Alkofer1994 ; ISHII1995 ; Zuckert1997 ; Chanfray_PRC2011 ; Carrillo2014 ; CARRILLO2016 ; Noro2024 ; Chanfray2024EPJA ; Chanfray2025 . These studies demonstrate that, at low energies, the structure of the effective meson Lagrangian is largely dominated by chiral symmetry. Bosonization of the NJL model yields an effective mesonic theory that reproduces the low-energy phenomenology of pseudoscalar (pion) and (axial-)vector mesons with reasonable accuracy. In addition, self-consistent soliton solutions of the bosonized NJL model have been extensively investigated to describe nucleons. Various approaches have been developed to describe these solitons and to connect nucleon properties with experimental data, including the emergence of chiral solitons from the bosonized NJL model Zuckert1994 ; Zuckert1995 ; Alkofer1995 ; Zuckert1997 , quark–diquark models Bentz2001 ; Carrillo2014 ; CARRILLO2016 , quark–meson descriptions, and structure function methods Weigel1997 ; Takyi2019 , as well as the appearance of $N$ and $N^{*}$ states in the parity-doublet model. Such soliton-based models provide a transparent mechanism by which baryons can be understood as solitons of effective meson theories rooted in an underlying quark-level description. These NJL chiral soliton solutions can be compared with the Skyrme type model by a gradient expansion in the limit of infinitely heavy (axial-) vector mesons Zuckert1995b . Further extensions have also addressed the role of strangeness, providing insights into hyperons in the context of NJL-based chiral solitons WEIGEL1994 ; Alkofer1995 .
@@
In-medium modifications of soliton properties, particularly in connection with the (fully or partial) restoration of chiral symmetry at finite density, have been investigated in several works, see Refs. CHRISTOV1990 ; Alkofer1991 ; Amore2000 . In-medium effects are implemented through density dependent values of the pion decay constant as well as the in-medium pion and sigma masses in a quark-meson theory in Ref. CHRISTOV1990 . It has ben suggested that the r.m.s charge radius of the soliton may increase by up to 17% near nuclear saturation density $\rho_{\rm sat}\sim 0.16\rm\ fm^{-3}$, and a phase transition from the chirally broken Goldstone phase to the chiral restored phase occurs near a baryon density of $\sim 3\rho_{\rm sat}$ CHRISTOV1990 .
@@
Another analysis suggests that the soliton solution vanishes near a critical quark chemical potential, indicating a phase transition where the bound soliton state dissolves to the constituents before the restored chiral symmetry Alkofer1991 . Furthermore, within a soliton bag model theory in the Wigner-Seitz approximation to find the quark wavefunction and bound sigma fields, it was predicted that an equilibrium among the nuclear and quark phase could exist around $\sim 6\rho_{\rm sat}$ and a complete quark plasma phase could be expected at $\sim 12\rho_{\rm sat}$ Birse1988 . Additionally, the soliton properties at finite temperature and density have also been examined, see Refs. Alkofer1991 ; Abu2012 ; Mao2013 ; Li2024 . The soliton description of baryons with vector mesons has also been employed to understand the properties of nuclear matter, such as incompressibility and binding energy, both in the Wigner-Seitz cell and the multi-skrymion approach. Since solitons are characterized by their r.m.s charge radii, this allows a self-consistent formalism of incorporating finite-size effects into nuclear and NS matter, linking the evolution of baryon size directly to the underlying soliton description Bentz2025 .
@@
As discussed in Ref. Fukushima2020 , the core and cloud picture provides a valuable framework to interpret baryonic structure at finite density. The outer, soft multi-pion cloud is directly linked to spontaneously broken chiral symmetry and the approximate Nambu–Goldstone nature of the pion. Its spatial extent is expected to grow with increasing density, consistent with the reduction of the in-medium pion decay constant. In contrast, the inner baryonic core is dominated by gluonic dynamics and is not directly tied to chiral symmetry, making it comparatively robust against moderate changes in density until the compact cores approach each other and begin to overlap. Well before this overlap sets in, the mesonic tails generate meson-exchange forces near saturation density ($\rho_{\rm sat}$). By further increasing the density, quark–antiquark pairs gradually delocalize, enhancing their mobility across multiple nucleons in a percolation-like process. This regime has been referred to as “soft” deconfinement. At sufficiently high densities, however, when the baryonic hard cores themselves overlap, quark matter is realized unambiguously. This transition is termed “hard deconfinement” and may be identified with the deconfinement of valence quarks, at which point the properties of dense quark matter are governed primarily by quark degrees of freedom. It can be quantitatively assessed from the internal nucleon hard core substructure  [^1] In Ref. Fukushima2020 , the discussed nucleon EoS corresponds to the de-confined quark matter (valid near the hard core overlap $\sim 8\rho_{\rm sat}$). However the EoS is represented as “nucleon EoS” as it is derived from the core of the nucleon..
@@
In the work presented in Ref. Fukushima2020 , the EoS for dense quark matter in the hard-deconfined phase is derived from the local nucleon EoS associated with the soliton’s hard core, constructed as a skyrmion of the pion field, evaluated under vacuum conditions. Although the restoration of chiral symmetry at high densities represents a natural physical expectation, its effects were not explicitly incorporated in the model. Instead, a qualitative argument was proposed to rescale the energy density and pressure using the in-medium pion decay constant as an approximation to account for chiral symmetry restoration. In contrast, the present study aims for a self-consistent approach to quantify the influence of chiral symmetry restoration on the density-dependent evolution of nucleon properties and the EoS. A central motivation of our framework is the description of the nucleon as a topological soliton emerging from the chiral dynamics of the NJL model, embedded within an in-medium effective background density-dependent scalar field that encapsulates the progressive restoration of chiral symmetry. As described below, the parameters of this NJL model will be adjusted to reproduce the skyrmion solution under the vacuum condition used in Ref. Fukushima2020 .
@@
To achieve the aim of this work, we leverage the path-integral bosonization approach of the underlying NJL model, as developed in Refs. Chanfray2010 ; Chanfray2025 . In summary, the NJL model is bosonized to express quark-level interactions in terms of mesonic degrees of freedom, facilitating the description of hadronic properties. The effective Lagrangian of the NJL model, after bosonization, includes scalar, vector, and pseudoscalar meson fields. The formalism establishes direct connections between the NJL parameters and the vacuum and in-medium modifications to key meson properties, including the pion mass, pion decay constant, vector coupling constant, and vector meson masses, all of which are evaluated in terms of the density-dependent scalar field. This formulation enables us to have a systematic accounting of the connection among the NJL model parameters and to both vacuum and in-medium nucleon properties, such as the soliton’s mass, radii, and EoS, while facilitating the exploration of scalar-field-induced or the restored chiral symmetry effects on mesons and solitons in dense nuclear matter.
@@
This paper is organized as follows. In Section. 2, the theoretical framework is provided, including the bosonized NJL model and the construction of the chiral soliton ansatz with vector mesons. In this section, details about the numerical implementation and its reliability by benchmarking it against existing results are presented as well. We discuss the sensitivity of the vacuum solution properties to the NJL model parameters and the effect of the chiral symmetry restoration on the nucleon structure and the resulting EoS in Section. 3. Finally, we summarise our findings and discuss the broader implications in Section. 4.
2 Theoretical Framework